OpenStudy (b2st):

I probably did something wrong... Help me what I did wrong? One cubic centimeter (1.0 cm3) of water has a mass of 1.0 multiplied by 10-3 kg at 25°C. Determine the mass of 1.5 m3 of water at 25°C. The answer is got was .15, however the answer was incorrect. What am I doing wrong? I'll put my whole process at the bottom.

OpenStudy (b2st):

\[1.5m ^{3} \times \frac{ (1cm^{3}) }{ 1\times 10^{-2} } = 150 cm^{3}\] \[150cm^{3} \times \frac{ 1.0 \times 10^{-3} kg }{ 1cm^{3}} = .15\]

OpenStudy (b2st):

^ That's how I did it, but it's still wrong.. When I entered the answer.

OpenStudy (anonymous):

Since you're cubing it, 1m^3 is not 1000cm^3 but rather 10^9 cm^3 so your answer would actually be 150000 kg

OpenStudy (b2st):

Okay I tried that before, and the site indicate it's incorrect.

OpenStudy (b2st):

wow never mind/ I'm so dumb, I forgot to input it in sig fig. lol